Operations Report – March 2nd

Crew 206 Operations Report 02 – MAR – 2019

SOL: 6

Name of person filing report: Gaspard Thieulin

Non-nominal systems: Generator

Notes on non-nominal systems: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.

Generator (hours run): 14,3 hours

Solar— SOC % – Turned off Gen : 69%, Turned on Gen : 54%

Diesel Reading – 75%

Propane Reading – 82%

Ethanol Free Gasoline – 7 gallons

Water (Auxillary tank) – Not in use

Gallons transferred: Not Applicable

Water in GreenHab – 257 gallons

Static to Loft Pump used – yes

Water Meter: 0141632,4 gallons

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Not used

Hours: 53,0 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Curiosity rover used: Not used

Hours: 77,3 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Not Used

Hours: 78,4 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Notes on rovers: Opportunity’s charger was down again (there might be an issue with water isolation somewhere as it happened on the two most wet days yet). Resetting the socket fixed it successfully.

ATV’s Used: N/A

Reason for use: Not used

Oil Added? N/A
ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

Hab Car used and why, where? Not used.

Staff Crew Car used and why, where? Off Campus

General notes and comments:

Summary of internet: Still on the fixing with BDM CapCom.

Summary of suits and radios: Nothing to report.

Summary of Hab operations: Cleaned the whole Hab this morning. We tried to put as many recipients for water leaks during today’s "storm". Also, many tarps seem to have moved because of the wind. We plan on a full check and reparation on the next EVA.

The line of electric sockets supporting the water heater stopped working without triggering any circuit breaker. Resetting the socket fixed it seemingly.

Summary of Green Hab operations: Nothing to report.

Summary of Science Dome operations: Nothing to report.

Summary of RAM operations: Nothing to report.

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: Nothing to report

Questions, concerns, and requests to Mission Support: Nothing more to report.